OpenStudy (ksaimouli):

slove exponential function 2.8^x=345

OpenStudy (ksaimouli):

\[2.8^x=345\]

OpenStudy (anonymous):

345 = 2.8^x log(345) = log(2.8^x) log(345) = xlog(2.8) x = log(345)/log(2.8)

OpenStudy (anonymous):

use calculator ;)

OpenStudy (ksaimouli):

no calc pro

OpenStudy (anonymous):

5,675

OpenStudy (anonymous):

x = 5,675 fine? understand?

OpenStudy (ksaimouli):

this problem we should not use calc